OpenStudy (anonymous):

Add. Simplify by removing factors of 1if possible 28zx + z-7x ----- ----- z^2-49x^2 z+7x

OpenStudy (dumbcow):

First factor bottom (z-7x)(z+7x) This shares the (z+7x) in other denominator so common denominator will be (z-7x)(z+7x) multiply 2nd fraction by (z-7x) on top and bottom \[=\frac{28xz +(z-7x)(z-7x)}{(z-7x)(z+7x)}\] expand the top and add like terms \[=\frac{z^{2} + 49x^{2}}{(z-7x)(z+7x)}\] top will not factor so this is answer